Last Tuesday, Cute Camel Woodcraft Company lost a portion of its planning and financial data when both its main and its backup servers crashed. The company's CFO remembers that the internal rate of return (IRR) of Project Gamma is 11.3%, but he can't recall how much Cute Camel originally invested in the project nor the project's net present value (NPV). However, he found a note that detailed the annual net cash flows expected to be generated by Project Gamma. They are: 


Year Cash Flow

Year 1  $2,400,000

Year 2  $4,500,000

Year 3  $4,500,000

Year 4 $4,500,000 


The CFO has asked you to compute Project Gamma's initial investment using the information currently available to you. He has offered the following suggestions and observations: 

  • A project's IRR represents the return the project would generate when its NPV is zero or the discounted value of its cash inflows equals the discounted value of its cash outflows-when the cash flows are discounted using the project's IRR. 

  • The level of risk exhibited by Project Gamma is the same as that exhibited by the company's average project, which means that Project Gamma's net cash flows can be discounted using Cute Camel's 8% WACC. 


Given the data and hints, Project Gamma's initial investment is _______ ,and its NPV is _______ (rounded to the nearest whole dollar). 

A project's IRR will _______ if the project's cash inflows increase, and everything else is unaffected.
Answer #1

1.
=2400000/1.113+4500000/1.113^2+4500000/1.113^3+4500000/1.113^4
=11985257.31214

2.
=-11985257.31214+2400000/1.08+4500000/1.08^2+4500000/1.08^3+4500000/1.08^4=974869.02362

3.
Increase